OpenStudy (anonymous):

For the geometric series 1 + 3 + 9 + 27 + . . . , find the sum of the first 10 terms A. 19,683 B. 29,523 C. 29,524 D. 88,573

OpenStudy (anonymous):

u can see the number increase by X3 so its going to be 1+3+9+27+81+243+729+2187+6561+19683=? as u can see A is wrong its between 29000 ~30000 so answer has to be B or C and finally when u add all the number the answer will be 29524 which is C

OpenStudy (anonymous):

here a=1, r= 3/1=3 sum of a geometric series is given as \[S_n=\frac{a(r^n-1)}{r-1}=\frac{1(3^{10}-1)}{3-1}=\frac{(59049-1)}{2}=\frac{59048}{2}=29524\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

in the formula R is the number or increase with is X3 so R=3 and A is the first number it starts which is 1 and N is the number of which terms which is 10